Circular No. 7679

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2001dm IN NGC 749 B. A. Beutler, M. Modjaz and W. D. Li, University of California at

Berkeley, report the discovery by LOTOSS (cf. IAUC 7514) of an apparent supernova (mag ∼ 15.4) in unfiltered images taken with the Katzman Automatic Imaging Telescope on Aug. 7.5 and 8.5 UT. The new object is at α = 1^{h}55^{m}40^{s}.55, δ = −29^{o}55′20′′.7 (equinox 2000.0), which is 6′′.9 west and 0′′.4 north of the nucleus of NGC 749. A KAIT image of the field taken on 2000 Dec. 28.2 showed nothing at this position (limiting mag ∼ 18.0).

V1178 SCORPII S. Mazuk, C. Venturini, R. J. Rudy and D. K. Lynch, The Aerospace

Corporation; and R. C. Puetter, Center for Astrophysics and Space Science, University of California at San Diego, report 0.8–2.5-μm spectrophotometry using the Lick Observatory 3-m telescope and Aerospace Near Infrared Imaging Spectrograph on July 11.16 UT, ∼ 20 days past peak brightness: “The spectrum was very low excitation and showed broad linewidths of ∼ 2500 km/s. Many permitted recombination lines of C I and N I were present, probably indicating an overabundance of these elements. The strongest lines were the Lyβ-fluoresced O I lines at 0.8446 and 1.1287 μm. He I 1.0831-μm, the line that frequently dominates the mid- and late-time nova spectra, is only about one-third the strength of H I Pγ.”

COMET C/2001 A2 (LINEAR) M. Kidger, Instituto de Astrofisica de Canarias, reports: “Monitoring

of the inner coma by the Spanish Comet Observers Team reveals two small photometric events that are possibly related to the ejection of small fragments from the nucleus. Representative CCD R magnitudes, using a 10′′ aperture and the USNO A2.0 catalogue for reference stars: July 23.996 UT, 12.62 (R. Ferrando, Pla D’Arguines, 0.26-m Schmidt-Cassegrain); 24.889, 12.80 (P. Manteca, Begues, 0.26-m Schmidt-Cassegrain); 25.936, 12.4 (D. Rodriguez, Madrid, 0.20-m Schmidt-Cassegrain); 29.853, 12.79 (Manteca); 30.933, 12.92 (Manteca); 31.881, 12.85 (Manteca); Aug. 2.896, 12.95 (Manteca); 3.844, 13.1 (Manteca); 7.84, 13.4 (R. Ligustri, Latisana, Italy, 0.2-m reflector); 7.852, 13.4 (Ferrando). The lightcurve indicates events of amplitude 0.4 and 0.2 mag initiating at ∼ July 25.0 and 30.0, respectively. For comparison, the major event around July 12 (cf. IAUC 7659, 7676) had an amplitude of 1.5 mag (10′′ aperture). This indicates that, if due to fragmentation, these later events were probably caused by the separation of small, short-lived splinters that may not have been directly observable.”

2001 August 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Brian G. Marsden